Dr. Drew runs 175 meters East for 70 seconds, then walks back 10 meters West in 20 seconds. What is his speed in meters/second for the entire trip?
Answer #1

Total distance travelled, S=175+10=185\: m

Total time, t=70+20=90\: s

So, speed of Dr. Drew for the entire trip,

                      v=\frac{S}{t}=\frac{185}{90}=\mathbf{2.056\: m/s}